# Agent Skill Framework: Perspectives on the Potential of Small Language Models in Industrial Environments

### v1

@@ Abstract
$\spadesuit$$\spadesuit$footnotetext: These authors contributed equally to this work.Agent Skill framework, now widely and officially supported by major players such as GitHub Copilot, LangChain, and OpenAI, performs especially well with proprietary models by improving context engineering, reducing hallucinations, and boosting task accuracy. Based on these observations, an investigation is conducted to determine whether the Agent Skill paradigm provides similar benefits to small language models (SLMs). This question matters in industrial scenarios where continuous reliance on public APIs is infeasible due to data-security and budget constraints requirements, and where SLMs often show limited generalization in highly customized scenarios. This work introduces a formal mathematical definition of the Agent Skill process, followed by a systematic evaluation of language models of varying sizes across multiple use cases. The evaluation encompasses two open-source tasks and a real-world insurance claims data set. The results show that tiny models struggle with reliable skill selection, while moderately sized SLMs (approximately – parameters) benefit substantially from the Agent Skill approach. Moreover, code-specialized variants at around parameters achieve performance comparable to closed-source baselines while improving GPU efficiency. Collectively, these findings provide a comprehensive and nuanced characterization of the capabilities and constraints of the framework, while providing actionable insights for the effective deployment of Agent Skills in SLM-centered environments111More results: https://anonymous.4open.science/r/LabAgentSkill-C0E7.
@@ 1 Introduction
Agent Skills Concept Agent skills, initially introduced by a leading code-oriented large language model (LLM) unicorn222https://github.com/anthropics/skills, are widely regarded as an effective framework for agent-centric context engineering. This approach can be viewed as a well-designed “static cheat-sheet” in contrast to a “dynamic cheat-sheet” Suzgun et al. (2025), enabling LLMs to focus on salient information through progressive context management. This approach substantially optimizes both the length and quality of context windows, enabling LLMs and agents to markedly reduce hallucinations while improving their use of tools, documents, and external knowledge resources. Unlike retrieval-augmented generation (RAG) (Lewis et al., 2020) systems—which rigidly encode textual or visual data into vector-space databases—this design directly leverages the in-context learning (ICL) capabilities of LLMs (Dong et al., 2024a; Brown et al., 2020b) and emergent reasoning to dynamically select the most pertinent information, instructions, and contextual knowledge. Furthermore, its progressive disclosure–oriented design maintains a tightly bounded effective context lengthMei et al. (2025), which, in turn, substantially increases the success rates of skill and tool selection.
@@ 2.2 Agent Skills Flourish, Not in SLMs
This framework was first introduced in a blog post by Claude333https://claude.com/blog/skills, and since then, their design has been widely adopted by major players, including VSCode, OpenAI, LangChain, and others. Ye et al. conceptualizes CE as an evolvable skill, proposing a two-layer framework termed "Meta CE" to automatically rewrite and optimize skill descriptions. Li demonstrates that, for many reasoning tasks, constructing a single-agent system equipped with a skill library achieves comparable accuracy to multi-agent systems while reducing token consumption and latency by approximately half. Numerous open-source libraries for Agent Skills—such as DeepAgents444https://github.com/langchain-ai/deepagents and the Agent Skill Collections555https://github.com/heilcheng/awesome-agent-skills have grown rapidly Chen et al. (2026), while also sparking discussions on skill safety and permissions Liu et al.. When using Agent Skills, the default models are typically large-scale proprietary models that require API calls. In the proposed engineering practice, model selection proves crucial for the skill routing of SKILL.md, wherein small-sized models Li et al. (2025b) often exhibit suboptimal success rates and low performance Belcak et al. (2025); Li et al. (2025a). Furthermore, research on the feasibility of Agent Skill for small models remains limited, and there is a lack of quantitative evidence on deployment-level efficiency gains—such as VRAM footprint and end-to-end latency.
@@ Appendix B Additional Experimental Settings
To manage conversational history under the context-window constraint, a deterministic message-trimming policy was applied. Specifically, the first message (typically the system prompt) was always retained, and only the most recent or messages were kept depending on the parity of the total message count. Core instructions were preserved while the prompt length was bounded, thereby reducing truncation risk and stabilizing inference cost in longer interactions. For formatting and metadata extraction from SKILL.md, the publicly available codebase is primarily used.666https://github.com/agentskills/agentskills/tree/main/skills-ref

### v2

@@ Abstract
$\spadesuit$$\spadesuit$footnotetext: These authors contributed equally to this work.Agent Skill framework, now widely and officially supported by major players such as GitHub Copilot, LangChain, and OpenAI, performs especially well with proprietary models by improving context engineering, reducing hallucinations, and boosting task accuracy. Based on these observations, an investigation is conducted to determine whether the Agent Skill paradigm provides similar benefits to small language models (SLMs). This question matters in industrial scenarios where continuous reliance on public APIs is infeasible due to data-security and budget constraints requirements, and where SLMs often show limited generalization in highly customized scenarios. This work introduces a formal mathematical definition of the Agent Skill process, followed by a systematic evaluation of language models of varying sizes across multiple use cases. The evaluation encompasses two open-source tasks and a real-world insurance claims data set. The results show that tiny models struggle with reliable skill selection, while moderately sized SLMs (approximately – parameters) benefit substantially from the Agent Skill approach. Moreover, code-specialized variants at around parameters achieve performance comparable to closed-source baselines while improving GPU efficiency. Collectively, these findings provide a comprehensive and nuanced characterization of the capabilities and constraints of the framework, while providing actionable insights for the effective deployment of Agent Skills in SLM-centered environments.
@@ 1 Introduction
Agent Skills Concept Agent skills, initially introduced by a leading code-oriented large language model (LLM) unicorn111https://github.com/anthropics/skills, are widely regarded as an effective framework for agent-centric context engineering. This approach can be viewed as a well-designed “static cheat-sheet” in contrast to a “dynamic cheat-sheet” Suzgun et al. (2025), enabling LLMs to focus on salient information through progressive context management. This approach substantially optimizes both the length and quality of context windows, enabling LLMs and agents to markedly reduce hallucinations while improving their use of tools, documents, and external knowledge resources. Unlike retrieval-augmented generation (RAG) (Lewis et al., 2020) systems—which rigidly encode textual or visual data into vector-space databases—this design directly leverages the in-context learning (ICL) capabilities of LLMs (Dong et al., 2024a; Brown et al., 2020b) and emergent reasoning to dynamically select the most pertinent information, instructions, and contextual knowledge. Furthermore, its progressive disclosure–oriented design maintains a tightly bounded effective context lengthMei et al. (2025), which, in turn, substantially increases the success rates of skill and tool selection.
@@ 2.2 Agent Skills Flourish, Not in SLMs
This framework was first introduced in a blog post by Claude222https://claude.com/blog/skills, and since then, their design has been widely adopted by major players, including VSCode, OpenAI, LangChain, and others. Ye et al. conceptualizes CE as an evolvable skill, proposing a two-layer framework termed "Meta CE" to automatically rewrite and optimize skill descriptions. Li demonstrates that, for many reasoning tasks, constructing a single-agent system equipped with a skill library achieves comparable accuracy to multi-agent systems while reducing token consumption and latency by approximately half. Numerous open-source libraries for Agent Skills—such as DeepAgents333https://github.com/langchain-ai/deepagents and the Agent Skill Collections444https://github.com/heilcheng/awesome-agent-skills have grown rapidly Chen et al. (2026), while also sparking discussions on skill safety and permissions Liu et al.. When using Agent Skills, the default models are typically large-scale proprietary models that require API calls. In the proposed engineering practice, model selection proves crucial for the skill routing of SKILL.md, wherein small-sized models Li et al. (2025b) often exhibit suboptimal success rates and low performance Belcak et al. (2025); Li et al. (2025a). Furthermore, research on the feasibility of Agent Skill for small models remains limited, and there is a lack of quantitative evidence on deployment-level efficiency gains—such as VRAM footprint and end-to-end latency.
@@ Appendix B Additional Experimental Settings
To manage conversational history under the context-window constraint, a deterministic message-trimming policy was applied. Specifically, the first message (typically the system prompt) was always retained, and only the most recent or messages were kept depending on the parity of the total message count. Core instructions were preserved while the prompt length was bounded, thereby reducing truncation risk and stabilizing inference cost in longer interactions. For formatting and metadata extraction from SKILL.md, the publicly available codebase is primarily used.555https://github.com/agentskills/agentskills/tree/main/skills-ref